OpenStudy (anonymous):

WILL FAN AND MEDAL SOMEONE TO CHECK MY ANSWER! A cookie company packages its cookies in rectangular prism boxes designed with square bases that have both a length and a width of 4in. less than the height of the box. Which polynomial below represents the volume of a box with height x inches? Question 4 options: V=x(x+4)2 V=x(x−4)2 V=x(x−4) V=2x(x+4) I think its option D?

OpenStudy (anonymous):

@jim_thompson5910 could you check my answer? ill medal

jimthompson5910 (jim_thompson5910):

x is the height, so x-4 is the length and also the width since `length and a width of 4in. less than the height`

jimthompson5910 (jim_thompson5910):

So L = x-4, W = x-4, H = x Volume = length * width * height V = L * W * H V = (x-4)*(x-4)*x V = ???

OpenStudy (anonymous):

x(x-4)2?

jimthompson5910 (jim_thompson5910):

I think you meant to say \(\Large x(x-4)^2\) right?

OpenStudy (anonymous):

yeah!:)

jimthompson5910 (jim_thompson5910):

it is correct

OpenStudy (anonymous):

thansk!

jimthompson5910 (jim_thompson5910):

no problem